A car radiator has a 6-liter capacity. If the liquid in the radiators 40% antifreeze, how much liquid must be replaced with pure anitfreeze to bring the mixture up to a 50% solutions? If I told you the answer is 1 liter, in your own words, using complete sentences explain how you would solve this problem





Answer by stanbon(57424)

A car radiator has a 6-liter capacity. If the liquid in the radiators 40% antifreeze, how much liquid must be replaced with pure anitfreeze to bring the mixture up to a 50% solutions? If I told you the answer is 1 liter,
--------------------
Equation:
acid -acid + acid = acid
----
0.40*6 - 0.40x + 1.00x = 0.50(6)
-----
Multiply thru 100 to get:
40*6 - 40x + 100x = 50*6
60x = 10*6
x = (1/6)*6
---
x = 1 liter (amt. of 40% removed and replaced with 100% antifreeze)
==========================
Cheers,
Stan H.